The winery cost accountants have determined that the wine is produced at an average cost of $20 per bottle or $240 per case. The distributor and retail margins average 35% 14%, and 35% respectively. Find the wines selling prices per bottle at each channel level.

1. Sales from the winery to distributers.

2. Sales by the distributers to retailers.

3. Sales by the retailers to the consumers.
Answer #1

Calculation of selling prices per bottle at each channel level:-

Unless otherwise stated in the problem margins are considered as sale margin.

Sale margin%= profit per unit/selling price per unit

Profit = selling price×sale margin%

Let selling price be the x

X= cost + profit

X = cost + x *sale margin%

x(1-sale margin%)= cost

X=cost/(1-sale margin%)

Selling price = cost per unit/(1-sale margin per unit)

1) sale from the winery to distributers

Selling price (sp)= 20/(1-0.35)

Sellinf price per bottle=30.76

2)sales by the distributers to retailers

Cost of the distributers is selling price of the manufacturer

SP=30.76/(1-0.14)

SP=35.77

3) sales by the retailers to the consumers

Cost of the retailers is selling of price of the distributers

SP= 35.77/(1-35%)

SP per bottle = 55.031